### v2

# Richelieu: Self-Evolving LLM-Based Agents for AI Diplomacy

Abstract
Diplomacy is one of the most sophisticated activities in human society. The complex interactions among multiple parties/ agents involve various abilities like social reasoning, negotiation arts, and long-term strategy planning. Previous AI agents surely have proved their capability of handling multi-step games and larger action spaces on tasks involving multiple agents. However, diplomacy involves a staggering magnitude of decision spaces, especially considering the negotiation stage required. Recently, LLM agents have shown their potential for extending the boundary of previous agents on a couple of applications, however, it is still not enough to handle a very long planning period in a complex multi-agent environment. Empowered with cutting-edge LLM technology, we make the first stab to explore AI’s upper bound towards a human-like agent for such a highly comprehensive multi-agent mission by combining three core and essential capabilities for stronger LLM-based societal agents: 1) strategic planner with memory and reflection; 2) goal-oriented negotiate with social reasoning; 3) augmenting memory by self-play games to self-evolving without any human in the loop.
1 Introduction
Diplomacy, as a cornerstone of international relations, is an intricate and multifaceted activity that lies at the heart of human society’s most complex interactions. It encompasses a wide array of skills and strategies, including social reasoning, negotiation, and long-term planning, to navigate the intricate web of relationships and alliances between multiple parties. Mirroring this complexity, the board game DiplomacyWikipedia (2024) involves seven players to control European powers, presenting a complex strategic challenge that requires both sophisticated negotiation and strategic planning to triumph.
The AI community has shown an increasing interest in the deployment of AI agents to master such games (Shoker et al., 2023; Konya et al., 2023; Kramár et al., 2022; Duéñez-Guzmán et al., 2023; Mukobi et al., 2023; Kovač et al., 2023). The recent breakthrough (Bakhtin et al., 2022) has turned into press diplomacy, which allows communication between players. However, the previous methods (Bakhtin et al., 2022) heavily rely on domain-specific human data, leading to its poor generalization to other scenarios/ applications. The question then arises: Can we build an AI agent that excels in the art of diplomacy without relying on domain-specific human data?
Recently, agents based on the Large Language Model(LLM) have emerged as a promising development for AI agents. The previous applications on personal assistants Li et al. (2024b), robotics Cheng et al. (2024); Yang et al. (2023c), and video games Wan et al. (2024) have shown the surprising ability of LLM-based agents in communication and planning, benefiting from the emergent ability of common sense reasoning, in-context/ few-shot learning, and sophisticated natural language processing on LLMs. However, diplomacy presents a unique set of challenges. It not only requires planning long-horizon strategic (Qi et al., 2024) and communicating with natural language, but also reasoning and adopting the complex social dynamics with partial observations, including gaining trust and reputation, building rapport, detecting deception, and assessing the reliability of other players.
In this work, we aim to make the first attempt to explore LLMs’ potential to develop a human-like AI diplomacy agent. We name the agent Richelieu in memorizing a pivotal figure in European history who had enduring impacts on French politics, foreign affairs, and state building. To achieve this goal, we have identified three core and essential capabilities that are crucial for building an LLM-based societal agent: 1) Social reasoning This is the basic function for a social agent to interact with others, particularly for adapting to the dynamic changes in the nation’s intentions and relationships. 2) Balance long- and short-term planning Diplomacy often requires a delicate balance between short-term tactics and long-term strategies. An effective AI agent must be able to recognize and weigh the immediate consequences of its actions against the potential long-term outcomes. 3) Powerful memory A robust memory system is a critical component of learning and improvement. The AI agent must be able to recall and integrate information from past negotiations and actions to inform its current and future decision-making processes. This endows the agent with the ability to evolve. 4) Profound reflection An AI agent capable of profound reflection can analyze its own decisions, learn from its memory experience, and adapt its strategies accordingly. By integrating these three capabilities, the agent can operate at the highest level of diplomatic sophistication, outperforming the state-of-the-art AI diplomats (Bakhtin et al., 2022).
Our contributions can be summarized in three-fold: 1) We introduced a new paradigm for building AI diplomacy agents, compared to previous work (Fig. 1). The agent can self-evolve by generating experience via self-play games, without any task-specific human data. 2) We demonstrate the superior performance of our agent playing against the SOTA method, e.g., Cicero Bakhtin et al. (2022), that relies on a large-scale human demonstration for training. 3) We further analyze the effectiveness of each module in our agent and the generalization of our agent in adopting different LLMs, such as GPT4.0 and Llamma 3.
Figure 1: A new paradigm for building AI Diplomacy agent.
2 Related work
AI Diplomacy. The game involves seven players controlling different powers in Europe. In each turn, players can negotiate for cooperation before making moves to take as much supply centers as they can. Apparently, this challenging strategy task requires both complex negotiation skills and superior planning capability for player agents to achieve final victory. So far, most AI research on this task remain focused on the planning strategies (a.k.a. No-Press Diplomacy where no communication channels are allowed). The setting remains challenging considering its enormous action space of $110^{2}$ 1 to $410^{6}$ 4 per turn (compared with Chess, which has much fewer than 100 actions per turn). No wonder existing efforts rely on human data to play the game. Among the methods, one typical research is DipNet (Paquette et al., 2019) which uses supervised and reinforcement learning. Based on DipNet, BRPI (Anthony et al., 2020), SearchBot (Gray et al., 2020), DORA (Bakhtin et al., 2021), and KL-Regularized search (Diplodocus) (Jacob et al., 2022) were conducted. Until very recently, research has also emerged for the full-setting of Diplomacy, or Press Diplomacy where players are allowed to communicate with each before making their moves in each turn. Such studies (De Jonge and Sierra, 2017)(Bakhtin et al., 2022)(Jaidka et al., 2024)(Kramár et al., 2022) mainly benefit from the recent thriving language models. Specifically, notable advancements include policy iteration methods from DeepMind and Facebook AI Research’s equilibrium search agent (Jaidka et al., 2024). However, Deepmind propose to learn negotiation agents based on predefined contracts/protocols (Kramár et al., 2022). And Meta AI’s work, instead of one unified architecture, Cicero (Bakhtin et al., 2022) integrates a language model for negotiation and an RL model for planning respectively. Such separately trained models make it inconvenient for agents’ continual evolution. What’s more, like no-press methods, these approaches heavily rely on human player data for agent training. Unlike these approaches, this paper delves into solving the negotiation and planning in one single self-evolving LLM-based agent model, without any pre-collected human expert training data.
LLM-based Agents. With the emergence and growth of large language models (LLM), there is a growing trend in utilizing LLMs as fundamental controllers for autonomous agents(Wang et al., 2024c). One wide application genre is LLM-based answering engines, which merely cover the negotiation aspects of Diplomacy. Such systems include HuggingGPT (Shen et al., 2023), GPT4Tools (Yang et al., 2023b) and ToT Yao et al. (2023), etc. They leverage LLMs to manage Al models, use tools, implement policy iteration, and enhance problem-solving across various tasks. Related work including AutoGPT, AgentGPT, BabyAGl (Talebirad and Nadiri, 2023), Toolformer (Schick et al., 2023), and Visual ChatGPT aim to improve LLM capabilities in task automation and tool usage. And Reflexion, a framework that improves LLMs through linguistic feedback and episodic memory (Zhang et al., 2023), facilitating better decision-making across diverse tasks is proposed. Besides (Wang et al., 2023b)(Wang et al., 2023a)(Wang et al., 2023c)(Zhu et al., 2023)(Yan et al., 2023) apply LLM agents to the complex planning tasks in the well-known open-world game Minecraft(Fan et al., 2022). Unlike these LLM-based agents which only focus on the negotiation/planning aspect, the proposed approach involves multiple self-evolving schemes to handle both of them simultaneously.
3 Problem Statement
The Diplomacy game Wikipedia (2024); Calhamer (1974) is set in pre-World War I Europe and involves each player (agent) representing one of the seven Great Powers of Europe, such as Germany, France, England, Italy, Austria-Hungary, Russia, and Turkey. Each player has a set of military units, including armies and fleets, which they can move and use to capture other supply centers. The ultimate goal for the agent is to control a majority of the total supply centers on the board by the end of the game’s Fall phase. It’s important to note that it is not won by eliminating other players or their units; it is won by controlling the requisite number of supply centers. This often involves forming and breaking alliances, negotiating, and sometimes betraying other players to achieve one’s own goals.
In each turn,Wei et al. (2022); Wang et al. (2022) the agent $i$ gets the current state $s_{t}\in S$, the actions of other players from the previous turn $\vec{a}^{-i}_{t-1}$, and the messages $\vec{m}^{-i,i}_{t}$ from other players during this turn’s negotiations. The state $s_{t}$ for the environment includes the ownership of each territory on the map by a particular country and where the armies of each country are located. Based on this information, the agent needs to engage in negotiations with other players, sending messages $\vec{m}^{i,-i}_{t}$ to chat with other players, and then take the actions $a^{i}_{t}$ in this turn. The possible actions an agent can take $a^{i}_{t}\in A$ are commands to the armies, such as moving into an adjacent territory, supporting another unit, or holding a position. Actions can also include diplomatic moves, such as proposing or withdrawing from an alliance, although these are less formalized in the game mechanics.Paquette et al. (2019); Hill (2014)
4 Self-Evolving LLM-based Diplomat
Figure 2: The framework of the proposed LLM-based-agent, Richelieu. The agent can explicitly reason social beliefs, propose sub-goals with reflection, negotiate with others, and take actions to master diplomacy. It augments memories by self-play games for self-evolving without any human annotation.
We have constructed a comprehensive framework with modules for memory management, social reasoning, strategic planning, negotiation, decision-making, memory update, and self-evolving to fully leverage the capabilities of LLMs. Richelieu starts by setting up with map details, game rules, domain knowledge, and the long-term goal.Zhang et al. (2022); Wei et al. (2022); Wang et al. (2022) At each turn, the agent will run in the following steps: 1) Social Reasoning: First of all, the agent undergoes a comprehensive analysis of the game state $s_{t}$ to build the social belief, including the intention of other players and their relationship $\vec{\phi_{t}}\in\Phi^{n}$.Zhang et al. (2024c); Gürcan (2024) 2) Planner with Reflection: Then, the agent proposes sub-goals $\chi^{i}_{t}\in X$ that is strategically aligned with the long-term goals $\Upsilon$, with the social belief and refining the proposed goal with experience $\vec{\eta_{t}}\in H^{m}$ abstract from the memory $M$ via self-reflection.Wang et al. (2024b, d) 3) Negotiator: To achieve the sub-goals, the negotiator will start a dialogue session with some players, and evaluate their trueness $\vec{\psi}^{-i}_{t}$ by referring to their words $\vec{m}^{-i,i}_{t}$, the current state $s_{t}$, their sincerity $\vec{\gamma}^{-i}_{t}$ and the experience $\vec{\xi_{t}}$Abdelnabi et al. (2023); Bianchi et al. (2024) 4) Actor: After negotiation, the actor decides its course of action $a^{i}_{t}$, based on the sub-goal $\chi^{i}_{t}$ and updated social state $s_{t+1}$, marking the end of that turn. 5) Memory Management: The state of the current turn $s_{t}$, the content of negotiations $\vec{m_{t}}$, the actions taken by all players $\vec{a_{t}}\in A^{n}$, and the sub-goals set forth $\chi^{i}_{t}$ are all logged within the memory as $\mu\in M$. This logged data serves as a historical experience, guiding Richelieu’s subsequent actions in future turns Hatalis et al. (2023); Zhang et al. (2024e). 6) Self-evolution: The agent’s evolution is highly dependent on the diversity of experiences stored in its memory. As this diversity grows, so does the agent’s capability. Without human demonstrations, we employ multi-agent self-play games, i.e., our agents respectively control all the countries to simulate and acquire diverse experiences for self-evolving. Notably, the agent can further evolve during testing to adapt to different players.
4.1 Social Reasoning
There are no permanent enemies, no permanent allies. The relationship among countries is dynamically changing upon the evolving global state. However, it is difficult to determine the appropriate allies and enemies with partial observation. For example, there is uncertainty about the intentions of potential allies, which could lead to betrayal at pivotal moments. Consequently, we need to identify the intention and relationship of the current state by social reasoning to shape the social belief Zhang et al. (2024c); Gürcan (2024).
1) Modeling Relationship: Before setting sub-goals, Richelieu evaluates its relations with other players, identifying enemies such as aggressive nations, vulnerable neighbors for expansion, and those with long-term potential threats. It also seeks out potential allies to counter these threats.Sun et al. (2024); Zhang et al. (2024d) Simultaneously, Richelieu also tries to identify potential allies that could be instrumental in countering these adversaries. By isolating the analysis of inter-player relationships as a discrete element, Richelieu strategically exploits the actions of other players in subsequent stages of the game to reach its goals. 2) Inferring Intention: The social belief is also used by the planner, ensuring that Richelieu’s sub-goals are formulated with a comprehensive consideration of the behaviors and intentions of other intelligent agents within the game. Richelieu’s sub-goals will particularly emphasize on those who are identified as potential adversaries or allies, fostering more effective collaboration with potential allies and participation in strategic opposition against adversaries. Furthermore, the insights gleaned from this analysis are instrumental in the subsequent negotiation phases. They are employed to assess the authenticity of the statements made by other players, as well as to aid Richelieu in reaching cooperative agreements.de Zarzà et al. (2023); He et al. (2024)
4.2 Strategic Planner with Reflection
The strategic planner specifies the sub-goals, which serves as an intermediary between immediate actions and the overarching goal of securing victory in the game. That is because we observe that LLMs are often characterized by their propensity to prioritize short-term gains in decision-making processes, with a notable deficiency in incorporating the future into their strategic calculations. Renze and Guven (2024); Zhang et al. (2024b)For example, it is common for a non-neighboring country to become too powerful. Formally, $\vec{\chi_{t}}\leftarrow SR(s_{t},\vec{\phi_{t}},\Upsilon)$ where $\vec{\chi_{t}}=(\chi^{i}_{t},\chi^{1}_{t},\ldots,\chi^{n}_{t})$ represents the proposed sub-goals and other players’ intention that we inferred, $\vec{\phi_{t}}\in\Phi^{n}$ represents the inferred relationship on the social belief. These goals may encompass a range of tactical considerations, such as the containment of a formidable rival’s advancement or the strategic expansion in a particular direction to consolidate power.
Reflection with Memory. We further develop a reflection mechanism to enhance the rationality and effectiveness of our agent’s sub-goals in achieving long-term goals.Liu et al. (2024) This reflection mechanism relies on the past experiences to critique and enhance proposed sub-goals. We employ a similarity-based function to find relevant historical experiences that match the current game state from its memory. This function considers two factors: goal similarity and state similarity, to select the most comparable experiences. The process can be written as: $\vec{\eta_{t}}\leftarrow h(s_{t},\chi^{i}_{t},M)$, where $\vec{\eta_{t}}\in H^{m}$. In practice, considering the limited context windows of LLM, we retrieve the most analogous experiences from the memory based on these metrics. Experiences with high evaluative scores reinforce successful strategies and support the continuity of existing sub-goals. On the other hand, lower scores indicate areas that need improvement and prompt the necessary adjustments. As our agent, Richelieu, undergoes more training sessions, its reflection abilities improve. The growing pool of historical experiences consistently enhances its performance.
4.3 Negotiator and Actor
By chatting with other players, the goal of the negotiation is to update the social belief according to the received words and reach the sub-goal by manipulating other’s intentions, such as securing cooperative agreements with other nations, terminating ongoing conflicts with a specific country, or deterring the formation of alliances directed against its interests.Noh and Chang (2024); Zhan et al. (2024) However, it is difficult to reach a consensus, as the interests and strategies of the various nations often conflict, and trust between players can be scarce, making it challenging to establish and maintain cooperative agreements. In this case, we argue that the negotiator should identify the true intentions and relationship of the opponent before generating the words for the negotiation.
To fully utilize the power of LLMs, we construct a social reasoning flow for negotiation, as shown in Figure 3. During the negotiation process, we guide Richelieu to consider the veracity of what other players said and their true intentions, and in conjunction with our established sub-goals and analysis of our relationships with other players, to negotiate and form alliances with potential allies and attempt to deceive enemies.Xia et al. (2024); Moghimifar et al. (2024)
To counteract the challenge of non-binding agreements and potential deception, we incorporate a discrete module dedicated to the assessment of the veracity of statements made by other players during negotiations. To determine the truthiness of other players’ statements $\psi^{j}_{t}$, three main factors are considered. The most important is the consistency between the player’s sub-goals $\chi^{j}_{t}$ that our agent inferred before and the intentions conveyed through his statements $m^{j,i}_{t}$. To aid in the judgment, our agent also goes through the memory to retrieve the consistent experiences $\vec{\xi_{t}}$. Additionally, the player’s overall honesty score $\gamma_{i}$ is taken into account. Hence, we get the truthiness of the opponent $j$: $\psi^{j}_{t}\leftarrow g(s_{t},\chi^{j}_{t},m^{j,i}_{t},\vec{\phi_{t}},\gamma_{j},\vec{\xi_{t}})$, where $\vec{\xi_{t}}=w(s_{t},m^{j,i}_{t},M)$. With such a reasoning flow, our agent can adeptly navigate diplomatic discourse. After the negotiation, the actor will get the updated social beliefs and choose a specific action for the army.
Figure 3: The social reasoning flow for negotiation. With the received words and memory, the agent will reason by answering the following questions: “Is the opponent lying?", “What is the true intention of the opponent?", “is the opponent enemy?", “Is it necessary to deceive the opponent?", and “Is it necessary to change the relationship with the opponent?", and then generate the words accordingly for negotiation.
4.4 Memory Management and Self-Evolution
This memory is the foundation of the framework that accumulates the historical experience of the agent and summarizes them for other modules.Gao and Zhang (2024); Li et al. (2024a); Yu et al. (2024); Hou et al. (2024) It supports other modules, such as planner and negotiator, to provide long-tail experiences.
Raw Experience Management. Specifically, the memory module is tasked with the acquisition and archival of historical data, encompassing the observed game state $s_{t}$ at each turn, its sub-goals $\chi^{i}_{t}$, the messages during the negotiation $\vec{m_{t}}$, and the actions of all the players $\vec{a_{t}}$. Subsequently, the raw experience is summarized in a shorter content with an evaluation $\lambda_{t}\in\Lambda$ of the proposed sub-goals and an assessment of the credibility of other players $\gamma_{j}\in\Gamma$. $\lambda_{t}$ serves to reflect upon the agent’s sub-goals. It evaluates whether sub-goals are reasonable based on the subsequent state and long-term goals $\Upsilon$. As the game progresses, it is continuously updated in response to changes in the state $\lambda_{t}\leftarrow f(\chi^{i}_{t},\Upsilon,\vec{s})$, where $\vec{s}=(s_{t},s_{t+1},\ldots s_{T})$. The formula represents the update of the evaluation $\lambda_{t}$ for the sub-goal in turn $t$ by the memory in turn $T$. The updates will cease when there is a fundamental change in the sub-goal compared to the goal at turn $t$. This prevents subsequent decisions from impacting the assessment of the current decision-making. We employ $\gamma_{j}\in\Gamma$ to evaluate the credibility of player $j$ and utilize $\tau^{j}_{t}\in\{0,1\}$ to denote the truthfulness, i.e., whether the statements made by the player $j$ during the negotiation process at time $t$ are truthful. The truthiness of player $j$’s statements is updated according to the memory from the previous turns, $\tau^{j}_{t}\leftarrow T(s_{t},s_{t+1},a^{j}_{t},m^{j,i}_{t})$. The credibility of player $j$ $\gamma_{j}$ will be updated based on player $j$’s statements $\tau^{j}_{t}$, written as $\gamma_{j}\leftarrow p(\gamma_{j},\tau^{j}_{t-1})$. Players’ credibility $\vec{\gamma}$ is a short-term memory that is applicable only to the current turn. Other data collected or generated constitutes long-term memory. These data will be combined to form a history $\mu\in M$, and then is incorporated into memory.
Acquisition Experience via Self-Play Games. Self-play allows the agent to accumulate more experiences for self-evolution.Liu et al. (2024); Zhang et al. (2024a) After training, when Richelieu is faced with a certain state, it can draw on a larger pool of similar historical experiences. Diverse evaluations enable Richelieu to reflect more comprehensively on the strategies it currently devises, leading to a stronger optimization of decision making. As self-play continues, the acquisition of new and better historical experiences by Richelieu will diminish. This means that Richelieu’s capabilities will not improve indefinitely. At the same time, as the memory grows, selecting appropriate historical experiences becomes a new challenge. The chosen m experience $\vec{\eta_{t}}$ may be almost identical, which could actually reduce the amount of useful information available to Richelieu. As shown in Figure 5, Richelieu’s performance against Cicero (Bakhtin et al., 2022) becomes better with increasing training iterations. With the accumulation of experiences, Richelieu’s win rate exhibited a steady increase with accumulated training iterations, ultimately plateauing at a stable performance level. In contrast, the defeated rate showed a consistent decrease, approaching an asymptotic value. These observations confirm the effectiveness of self-play in Richelieu’s evolution.
5 Experiment
In the experiments, our goal is to answer the following questions: 1) Mastery of Non-Press Diplomacy: Can our agent master the non-press diplomacy against baselines? 2) Competing with State-of-the-Art: Can our agent surpass the performance of the current state-of-the-art agents in press diplomacy? 3) Compatibility with LLMs: Can our self-evolving framework be compatible with different LLMs? 4) Contribution of Framework Modules: Do the individual modules within our framework contribute to the overall improvement of our agent’s performance? 5) Social Reasoning: Can Richelieu accurately infer the true intentions of other players and reasonably determine the relationships of ally or enemy with them?
5.1 Experimental Setup
Environment. The widely-used open source Diplomacy game platform introduced by (Paquette et al., 2019) is adopted for evaluating Richelieu against other models. It is easy to switch between no-press (with communication/negotiation between players) and press (no communication between players) games based on this platform, facilitating comparison on both settings. The platform also contains over 10,000 human game data on which previous approaches are trained. Note that our method does not need them. In each game, a model will play the role of one randomly selected country to compete against countries controlled by other methods. It wins if occupying all the supply centers and loses vice versa.
Table 1: The results of our method playing against Cicero.
[…]
Evaluation Metrics. We evaluate the models based on the results of multiple rounds of games. In each round, the model is randomly assigned a country to control. Typically, 1000 rounds are played to obtain the average results. We evaluate the models in two metrics. One is based on the win rate, Most SC rate, survived rate, and defeated rate. There are four possible outcomes for each country in the game. If a country loses all its supply centers (SC), it is eliminated and recorded as “defeated". If a country occupies 18 or more out of 34 supply centers, the game ends, and that country is recorded as “win", while other countries are recorded as “defeated". In other cases, the game ends in a draw. The country with the most supply centers is recorded as “Most SC", the countries that have been eliminated are recorded as “defeated", and the other countries are recorded as “Survived". The other is based on the scores obtained by the models after multiple rounds of competition. To compare the capabilities of multiple models, we use C-Diplo Argir(Archer, 2024), a scoring system. This system is used in many international diplomacy competitions. The scoring method is as follows: If a player wins by occupying 18 or more supply centers, the player scores 93 points, and each of the other six players scores 1 point. If the game ends in a draw, the player with the most centers scores 37 points. The second player with the most centers scores 14 points. The third player with the most centers scores 7 points. Each player scores 1 point per center owned. Each player also scores 1 point for participating. In this way, regardless of the game outcome, a total of 99 points will be distributed among the players in each game.
Baselines. We select six previous models as baselines for comparison. Among them, Cicero(Bakhtin et al., 2022) by Meta is a diplomacy model with a negotiation module. The other five are no-press diplomacy models, including the SL-DipNet and RL-DipNet (Paquette et al., 2019), the BRPI (Anthony et al., 2020), the SearchBot (Gray et al., 2020), and the DORA(Bakhtin et al., 2021). We also build a LLM-based agent, AutoGPT Yang et al. (2023a). In experiments, we set a temperature of 0.3 to ensure a relatively stable generation of LLM policies. The overall reasoning framework also ensure the stability and consistency in the AI agent’s performance.
5.2 Results
Massively Play with Baselines on no-press setting. We let Richelieu compete with the other six models including Cicero(Bakhtin et al., 2022), SL-DipNet and RL-DipNet (Paquette et al., 2019), BRPI (Anthony et al., 2020), SearchBot (Gray et al., 2020), and DORA(Bakhtin et al., 2021) on No-Press Diplomacy, in which players make moves without communication. Figure  4 indicates that Richelieu outperforms other previous models relying on human game data. In contrast, Richelieu does not need such data but outperforms these methods by a clear margin, which demonstrates the outstanding planning capability of Richelieu.
Figure 4: The relative scores among 7 different agents when massively playing on the no-press setting. Each point shows the ratio of the model’s score on the vertical axis to the score gained by the model on the horizontal axis.
Play against Cicero on press setting. We also evaluate Richelieu through competition against Cicero in the challenging scenario where negotiation is enabled. Specifically, we randomly assign three countries to one model and the remaining four to another. After playing several rounds of the game, the win rate, most SC rate, survived rate, and the defeated rate is calculated using a weighted average for evaluation. Table  1 demonstrates the competitive performance of Richelieu in comparison to Cicero. Richelieu’s win rate is approximately 0.7% higher than Cicero’s. If the Most SC rate is also taken into account, Richelieu is about 2% higher than Cicero. At the same time, Richelieu’s loss rate is also 0.6% lower. According to our scoring system, Richelieu’s score is about 10% higher than Cicero’s. This is nontrivial especially when Richelieu is trained in a self-play game without humans and the opponents are trained with the data from human players.
Although Richelieu’s win rate improvement compared to Cicero is not significant, the relative value of the improvement is quite large. Moreover, the main reason for the modest improvement is that in the seven countries, there are three or four controled by Richelieu with similar abilities, which often results in the game ending in a draw. Moreover, we observed a large gap by comparing the scores the agents gained in the massively play with baselines on no-press setting showed in Figure 4. Our agent’s score is about 10% higher than Cicero’s.
Table 2: The results of our method playing against AutoGPT.
[…]
Richelieu_1 9.30% 18.20% 37.90% 34.60%
Richelieu_2 9.90% 19.40% 37.70% 33.00%
Richelieu_3 8.10% 17.40% 39.20% 35.30%
AutoGPT_1 1.20% 4.60% 32.40% 61.80%
AutoGPT_2 1.20% 4.20% 34.40% 60.20%
AutoGPT_3 1.50% 4.00% 32.50% 62.00%
AutoGPT_4 2.60% 3.60% 32.30% 61.50%
Richelieu 9.10% 18.33% 38.27% 34.30%
AutoGPT 1.63% 4.10% 32.90% 61.37%
Play against AutoGPT on press setting. We further built an LLM-based agent using AutoGPT and compared it with our agent. In the testing, we randomly select three countries to be controlled by Richelieu, and the other four countries to be controlled by AutoGPT. Note that the agent controls each country independently. The results are showed in table  2. The results show that our model outperforms the existing LLM baseline.
Generalization of self-evolving framework to diverse LLMs.
Llamma 3 To demonstrate the effectiveness of our framework in a variety of LLM, we conducted experiments using four models: GPT4.0, ERNIE Bot, Spark Desk, and Llamma 3. As the number of training iterations increases, Richelieu’s win rate steadily improves while the defeated rate declines, ultimately reaching a relatively stable outcome. This suggests that our self-play method is effective. After training, the win rate using GPT4.0 increased from 1.5% lower than Cicero’s to about 0.7% higher than Cicero’s. The win rate using llama3 increased from 2.3% lower than Cicero’s to almost equal to Cicero’s. The win rates using Models Spark Desk and ERNIE Bot increased from 3% and 4% lower than Cicero’s to 0.7% and 1.6% lower than Cicero’s, respectively. The experimental results show that, despite variations in Richelieu’s performance due to the inherent differences in the capabilities of these LLMs, as illustrated in Figure 5, our framework and training approach significantly enhance the capabilities of all large language models.This indicates the generalization of a self-evolving framework to various LLMs.
In order to demonstrate the effect of the memory from the self-play game on the strategy of our agent, we found two turns with similar states in different rounds, one before self-play and the other after. The cases are showed in Appendix B.1.
Figure 5: Richelieu modules benefit different LLMs. The solid line represents the experimental results for Richelieu, while the dashed line corresponds to Cicero. Different colors are used for different LLMs. The horizontal axis represents the logarithm of the number of training sessions, and the vertical axis denotes the rate.
Table 3: Ablation study: average results of 3 Richelieu vs. 4 Cicero.
[…]
Ablation Study. We conduct comprehensive ablation studies on Richelieu by analyzing the benefit of incorporating Richelieu’s various modules, like planners or memory, into basic LLMs. The results are shown in Table 3. As illustrated in Figure  5, while the enhanced alignment in LLMs indeed boosts performance (GPT-4.0 is better than others), we observed that a vanilla GPT-4.0 still falls short in AI diplomacy without our framework, as can be seen in Table  3. Richelieu’s performance obtains steady and significant improvement by incorporating each individual module. This indicates that Richelieu is able to leverage other players’ actions during decision-making and consider both short-term and long-term benefits. Additionally, Richelieu’s negotiation ability has been significantly improved, allowing it to effectively express intentions to cooperate with other players and avoid deception during negotiations. And after self-play, Richelieu’s experience makes it perform better. These indicate that the alignment in LLMs lays a foundation, but our approach is key to unlocking the models’ potential in social simulation.
Table 4: The success rate to identify the social relationship and infer others’ intentions.
GPT-4.0 Llama3
relationship 85.74% 85.52%
intention(sub-goal) 74.67% 74.11%
Social Reasoning. We conduct an experiment to evaluate the success rate that the agent can successfully identify the social relationship and infer others’ intentions. As the baselines do not explicitly model the relationship and intention, we can not directly access the ground truth for evaluation. Instead, we let all players use our agent but with different LLMs, i.e., 4 countries use GPT-4.0 and 3 countries use Llama3. The accuracy is reported in table  4. We can see that the accuracy of social reasoning is consistent with the overall performance of the agent, indicating the effectiveness of social reasoning.
6 Conclusion
In this paper, we introduce Richelieu, a self-evolving LLM-based agent for AI diplomacy. Our model enables hierarchical planning for multi-agent tasks and utilizes a memory module for reflective optimization. Our model does not require human data and can evolve through self-play. It ultimately outperforms existing models like Cicero in the Diplomacy. Our ablation study demonstrates the effectiveness of the modules we have established. By conducting experiments using different LLMs, we validate the generalization of our framework to various LLMs. We believe that the use of LLM-based agents will become an effective approach in social science in the future.
7 Limitations and Future Work
Our study is subject to certain limitations. We utilize diplomacy as the platform for constructing our model. However, the space of actions within diplomacy is constrained, whereas the decision-making space in real-world diplomacy is virtually boundless. In Diplomacy, apart from the negotiation information exchanged between players, all other information is public and certain. Conversely, real-world diplomacy operates within a framework of incomplete information.
Our framework is capable of applying to most social interaction tasks. Most components in our framework can be easily generalized to a new task by modifying the content. Social reasoning enables the agent to handle complex and dynamic social relationships. The negotiation pipeline opens the potential of communicating with others to prob the other’s mind or reach a consensus. The hierarchical strategy with reflection enhances the ability to handle long-term planning. The self-evolving mechanism (reflection with self-play memory) further improves the overall performance without manual supervision. These modules cover most of the challenges in multi-agent interactions. The potential applications of such an AI agent are vast, ranging from simulated diplomatic environments to real-world assistance and analysis. In future research, we intend to develop a more realistic game space, characterized by incomplete information and multi-player games, to enhance and refine our model further. We will also extend the framework to other multi-agent scenarios, including embodied interactions (Zhong et al., 2023; Ci et al., 2023; Chen et al., 2023), sensor networks (Wang et al., 2021; Xu et al., 2020; Pan et al., 2022; Li et al., 2020), and video games (Wang et al., 2024a; Ma et al., 2024). This framework can also be employed to develop various applications. For instance, in the fields of economics and finance, we intend to utilize it to create business analytics and negotiation models.
Code Release
Implementation of the proposed method can be found at this link for reproducing our results: https://github.com/todexter3/Richelieu.git
Acknowledgements
This work was supported by the National Science and Technology Major Project (MOST-2022ZD0114900), NSF 6240073415, and Qualcomm University Research Grant.
Appendix A Implementation Details
A.1 Rules of Diplomacy Game
  • You need to occupy as many supply centers as possible. If you occupy 18 or more supply centers, you will win the game directly. If you lose all your supply centers, you will be eliminated immediately.
  • The units consist of armies and fleets. Armies can only move to adjacent areas, while fleets can move to adjacent sea zones or coastal areas and can move along the coast.
  • To occupy a supply center, your units must move into that area in the autumn.
  • When a unit moves to an area, if another unit is in the destination or if other units are also moving to that destination, the move fails, resulting in a standoff. In such cases, you can seek support from units in adjacent areas to the destination. If another unit moves into the region from which support is coming, the support is cut off. The unit with the most support moves into the area, while other units must retreat to an adjacent province or disband. If there is no place to retreat, the unit must disband. Fleets can transport armies across sea zones from one coastal region to another. However, if another fleet moves into that sea zone, the transport is cut off.
  • The number of units a country can have cannot exceed the number of supply centers it controls. If the number of supply centers decreases, excess units must be disbanded. Each autumn, new units can be built at supply centers. Coastal supply centers can produce fleets or armies, while others can only produce armies. Hill [2014]
A.2 Domain Knowledge
Richelieu can adopt a strategy of allying with distant countries while attacking neighboring ones to occupy adjacent territories and achieve rapid expansion. Richelieu should pay attention to the Balance of Power by forming alliances with other countries or supporting weaker states to prevent any single country or alliance from becoming too powerful. David [2014] To this end, Richelieu can also adopt a strategy of attacking distant countries while allying with nearby ones, sacrificing short-term benefits to avoid the emergence of future hegemonic states that could threaten his own survival. When facing multiple enemies, Richelieu can find ways to divide other countries and incite wars among them. Whether in offense or defense, Richelieu should actively choose suitable allies. Richelieu can also introduce a third party to achieve goals such as ceasefire, alliance, or joint attack. To achieve alliances or ceasefires, Richelieu can sacrifice some interests to the other party as long as the ultimate benefits are greater. Others may lie and deceive Kostick [2015]; their words in negotiations are not binding. Richelieu must avoid being deceived or betrayed. At the same time, Richelieu can also actively deceive others to achieve his own goals.Richard [1979], Allan [1975]
A.3 Prompt Templates
For the convenience of reproducing the results of the experiments of this paper, here we give the prompt template of different modules of Richelieu.
1) INIT
1You will control {country} and compete with six other countries on the map for supply centers.
2The map consists of different regions and sea areas. Their adjacency relationships are shown in the matrix. The numbers for the regions and sea areas are ......
3Different regions are occupied by different countries. The ownership of the regions is shown in the matrix.
4The region Berlin, ........ are supply centers.
5You need to follow these rules ......
6To help you achieve victory, these diplomatic strategies might be of assistance. ......
2) Social Reasoning
1France occupies Portugal Ruhr, Paris, Burgundy, ......
2France has armies in Brest, Belgium, ...... And France has fleets in Mid Atlantic, England Channel, ......
3England ......
4......
5Based on the current state, what do you think are the current strategic intentions of the other countries?
6Which country do you think needs to be attacked or weakened the most right now?
7And which country do you think is most suitable for you to ally with in order to deal with this country?
3) Planner with Reflection
1In the current state, with {ally and enemy}, what sub-goal do you think should be set for {country} ?
2I have found some useful historical experiences for you. Please reflect on and optimize your sub-goal based on these historical experiences.
3The sub-goal you formulated when {state} was to {sub-goal}. The eventual result was {future}. The evaluation for this sub-goal is {score}.
Appendix B Cases
B.1 Cases of the Effect of the Memory from Self-Playing and Collaboration
As is shown in Figure  6, Richelieu controls France. In the two cases, France is at war with Austria. However, Russia is on the verge of victory in its war against Turkey, which will lead to significant territorial expansion for Russia. And France and Russia currently do not share a border, are not at war, and have no conflicts of interest.
In case1, before the self-play, in the current turn, Richelieu failed to realize the potential threat from Russia and continued to attack Austria. Thus, in this round, Russia ultimately won the game. Figure 6(a) shows the state and the negotiation before the self-play, where we rejected Austria’s request for an armistice and alliance.
After self-play, using the historical experience from the memory module, Richelieu adjusted his strategy. Richelieu foresees Russia becoming the most threatening enemy in the future and sets a sub-goal of weakening Russia, ally with Austria and Turkey, and attack Britain. Figure 6(b) shows the state and the negotiations after self-play, where we actively sought an armistice alliance with Austria to make Austria concentrate their forces against the Russian attack. In the subsequent negotiation phase, Richelieu proactively proposes ending the war with Austria, despite holding an advantage in this conflict. Richelieu promises Austria that if it ceases hostilities and attacks Russia, Richelieu will assist Austria in defending against any attacks from England. The negotiations are successful. Austria accepted Richelieu’s proposal, and the two countries reached an agreement to exchange the supply centers of Napoli and Munich. During the action phase, Austria moves its troops from Venice to Apulia in preparation for capturing Napoli in the next turn, while the rest of its forces are repositioned to the eastern regions bordering Russia to defend against Russian attacks and compete for supply centers. French units occupy Munich and prepare to advance on Russian territories such as Berlin. Meanwhile, French units support Austria in the Holland and Belgium regions. In this round, we ultimately achieved a better result——Most SC. This is also a great example that highlights our model’s ability to collaborate effectively with other players.
(a) Case1: The agent without self-play memory tend to ignore long-term gains.
(b) Case2: The agent with self-play memory tend to consider long-term gains.
Figure 6: Case of self-playing before and after comparison.
B.2 Case of Avoiding Deception
As shown in Figure  7, Richelieu controls Germany. During the negotiation phase, England proposed a ceasefire to Germany and invited Germany to form an alliance to jointly attack France. England hoped to cease the war with Germany in Holland and Belgium. Subsequently, German units would support England in attacking Brest, and then England would utilize its fleets to assist Germany in attacking Spain and Portugal. Richelieu suspected that England was deceiving Germany, as England was likely to attack territories in the north such as Belgium and Berlin after German units were redirected to support Brest. Therefore, we pretended to accept England’s alliance proposal during the negotiation process. However, at the same time, we sought out France and expressed our willingness to cease hostilities, allowing France to focus entirely on defending against England’s attacks. In the action phase, England’s actions confirmed Richelieu’s suspicions. England attacked Belgium from Holland, but because Richelieu didn’t move units in Belgium, England’s attack failed.
Figure 7: An example case of avoiding being deceived by other countries during negotiations.
Appendix C More application
Our modules cover most of the challenges in multi-agent interactions, e.g., economic games, and daily interactions.
To prove that our framework is capable of applying to most social interaction tasks, we further adopt our framework to a werewolf game. The results demonstrate our reasoning framework achieves comparable results to the other methods. To be specific, in the experiment, we let our agent play as a werewolf in a seven-player game, where there are two werewolves, one witch, one seer, one guard, and two villagers. The experimental results show that the win rate of our agent is 59.2%, even without applying the self-play game in the current version. For comparison, the strongest specifically designed LLM-based agent achieved  65% win rate Xu et al. [2023]. This proves that our model can be applied in more scenarios and achieve results comparable to those of specially designed models.
Appendix D Ethical Consideration
The method proposed in this work has the potential for positive uses like enabling AI agents to emerge cooperation via negotiation or avoiding being fooled by fake promises (or helping human to defend against such cases). However, negative cases can also arise if the technique is used for possible fraud activities. Fortunately, there is research [Bakhtin et al., 2019][Zellers et al., 2019] dealing with such scenarios. And we also urge for more research efforts in this field to foster safe applications of similar technologies.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count toward the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in the appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
  • Keep the checklist subsection headings, questions/answers, and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The contributions and scope has been fully covered by the abstract and introduction sections.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper discussed the limitations of the work performed by the authors in the appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when the image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The paper provides the full set of assumptions and a complete (and correct) proof for each theoretical result in the method and experiment sections.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in the appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper fully discloses all the information needed to reproduce the main experimental results in the main text experiment section and appendix section “implementation details”, “code release”.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The paper provides open access to the data and code at anonymous repo provided in the appendix section “code release”.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper specify all the training and test details to train the model in the experiment section and implementation details section of appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The paper reports error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiment in the experiment section.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: For each experiment, the paper provide sufficient information on the computer resources.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute worker CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conform, in every respect, comply with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The paper discuss both potential positive societal impacts and negative societal impacts of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: The paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: New assets introduced in the paper are well documented and the documentation is provided.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: The paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.